# Sequential FDR and pFDR control under arbitrary dependence, with application to pharmacovigilance database monitoring

### v1

@@ Abstract
We propose sequential multiple testing procedures which control the false discover rate (FDR) or the positive false discovery rate (pFDR) under arbitrary dependence between the data streams. This is accomplished by “optimizing” an upper bound on these error metrics for a class of step down sequential testing procedures. Both open-ended and truncated versions of these sequential procedures are given, both being able to control both the type I multiple testing metric (FDR or pFDR) at specified levels, and the former being able to control both the type I and type II (e.g., FDR and the false nondiscovery rate, FNR). In simulation studies, these procedures provide 45-65% savings in average sample size over their fixed-sample competitors. We illustrate our procedures on drug data from the United Kingdom’s Yellow Card Pharmacovigilance Database.
@@ 1 Introduction
Hart and Weiss, (1997) used linear programming to find a sharp upper bound on the probability of rejection for a class of multiple testing procedures in terms of the true, marginal c.d.f.s of the $p$-values. Calculating their bound therefore requires knowledge of at leas the marginals of the true joint distribution. Guo and Rao, (2008) also used optimization but to find an upper bound on the false discovery rate (FDR) in terms of a step down procedure’s step values, that holds under arbitrary joint distribution of the $p$-values. By choosing the step values to control the upper bound, they produce a step down procedure whose FDR control holds under arbitrary joint distributions. The current paper extends this line of research in the following ways. First, by extending to the sequential setting an FDR controlling procedure without requiring or imposing assumptions on the joint distribution of the $p$-values; the open-ended versions of these sequential procedures allow the possibility of simultaneously controlling both type I and type II error metrics (e.g., FDR and the false nondiscovery rate, FNR). Secondly, by removing an implicit assumption on the $p$-values in Guo and Rao’s proof that is stronger than validity of $p$-values and applies to both the sequential and fixed-sample settings; see Appendix A for details. Finally, by extending this methodology to positive false discovery rate (pFDR) and its type II analog.
@@ 2.1 Data Streams, Hypotheses, and Error Metrics
In order to implement our proposed procedures, the marginal type I and II error probabilities of component test statistics on each stream will need to be controlled controlled at certain levels; see (4)-(5) below. Beyond that, we make no assumptions about the dimension of the sequentially-observed data $X_{n}^{(j)}$, which may themselves be vectors of varying size, nor about the dependence structure of within-stream data $X_{n}^{(j)},X_{m}^{(j)}$ or between-stream data $X_{n}^{(j)},X_{m}^{(j^{\prime})}$ ($j\neq j^{\prime}$). In particular there can be arbitrary “overlap” between data streams, an extreme case being that all the data streams are the same, which is equivalent to testing multiple hypotheses about a single data source. For any positive integer $j$ let $[j]=\{1,\ldots,j\}$. For each data stream $j\in[J]$, assume that there is a parameter vector $\theta^{(j)}\in\Theta^{(j)}$ determining that distribution of the stream $X_{1}^{(j)},X_{2}^{(j)},\ldots$, and it is desired to test a null hypothesis $H^{(j)}$ versus the alternative hypothesis $G^{(j)}$, where $H^{(j)}$ and $G^{(j)}$ are disjoint subsets of the parameter space $\Theta^{(j)}$ containing $\theta^{(j)}$. The null $H^{(j)}$ is considered true if $\theta^{(j)}\in H^{(j)}$, and false if $\theta^{(j)}\in G^{(j)}$. The global parameter $\theta=(\theta^{(1)},\ldots,\theta^{(J)})$ is the concatenation of the individual parameters and is contained in the global parameter space $\Theta=\Theta^{(1)}\times\cdots\times\Theta^{(J)}$.
@@
The general notation (1) includes fully-sequential sampling where the streamwise sample sizes may take any value $1,2,\ldots$ ad infinitum, but other set ups including group sequential, truncated, and even fixed sample size can be included in this as well. For example, the $n$th “observation” $X_{n}^{(j)}$ in the $j$th stream may actually be the $n$th group $X_{n}^{(j)}=(X_{n,1}^{(j)},\ldots,X_{n,\ell}^{(j)})$ of size $\ell$. Moreover, the group size $\ell$ may vary with $n$ and may even be data-dependent, e.g., determined by some type of adaptive sampling. Similarly, truncated sequential (or group sequential) sampling can be implemented for the $j$th stream by defining $X_{n}^{(j)}=\emptyset$ for all $n>\overline{N}^{(j)}$ for some stream-specific truncation point $\overline{N}^{(j)}$, or globally for all streams by replacing statements like “for some $n$” in what follows with “for some $n\leq\overline{N}$,” for some global truncation point $\overline{N}$. One may represent a fixed sample size in the $j$ the stream in this way by taking $\overline{N}^{(j)}=1$, or in all streams with $\overline{N}=1$.
@@
For any multiple testing procedure under consideration, let $V$ denote the number of true null hypotheses it rejects (i.e., the number of false positives), $W$ the number of false null hypotheses it accepts (i.e., the number of false negatives), and $R$ the number of null hypotheses it rejects. The number of null hypotheses accepted is therefore $J-R$. Under true value of the parameter $\theta$, the false discovery and non-discovery rates (Benjamini and Hochberg,, 1995, FDR, FNR) are
@@
where $x\vee y=\max\{x,y\}$. Similarly, the positive false discovery rate pFDR (Storey,, 2002) and its type II analog the positive false nondiscovery rate pFNR are defined as
@@ 2.2 Test Statistics and Critical Values
The theorems in this paper will give upper bounds on the FDR and pFDR, and their type II, of versions sequential procedures implemented with arbitrary step values $\bm{\alpha},\bm{\beta}$. In practice we recommend using some of the commonly used values such as those recommended by Benjamini and Hochberg, (1995) or Benjamini and Liu, (1999). For desired value $q_{1}\in(0,1)$ of the type I error metric (FDR or pFDR), the former are
@@
Here $x\wedge y=\min\{x,y\}$. We recommend the using the same expressions for $\bm{\beta}=\bm{\beta}(q_{2})$ for desired value $q_{2}\in(0,1)$ of the type II error metric (FNR or pFNR). In our examples below we will refer to (12) and (13) as the BH and BL step values, respectively.
@@ 2.3 Simple hypotheses and Wald approximations
In order to apply the above to the $j$ the stream in the multiple testing set up for given step value vectors $\bm{\alpha}$ and $\bm{\beta}$, a slight extension of Bartroff and Song, (2020, Theorem 5.1) shows that choosing
@@
satisfies (4)-(5) up to Wald’s approximations, and that $\widetilde{\alpha}_{k}+\beta_{k}\leq 1$ and $\alpha_{k}+\widetilde{\beta}_{k}\leq 1$ for all $k\in[J]$. Although, strictly speaking, the inequalities (4)-(5) will only be guaranteed to hold approximately using these approximations, Hoel et al., (1971) show that the actual type I and II error probabilities can only exceed the nominal values by a negligibly small amount in the worst case, and the difference approaches 00 for small nominal values, which is relevant in the present multiple testing situation where we will utilize fractions of $\alpha$ and $\beta$. Alternatives to the Wald approximations in the simple vs. simple set up using a Monte Carlo or to replace the terms in (16) by the conservative values $\log\beta$ and $\log(1/\alpha)$, respectively, for which (14)-(15) always hold (see Hoel et al.,, 1971) and proceed similarly.
@@ 4.3 Truncated, rejective procedures controlling FDR or pFDR
    regardless of the dependence between data streams, if $\gamma_{1}>0$ satisfy (25) for this procedure.
@@ 5 Simulation studies
In this section we present the results of simulation studies of the sequential step down procedures described above in Sections 3 and 4 in order to evaluate their operating characteristics and compare them to analogous fixed-sample procedures. All computations in this section and the analysis of the UK Yellow Card pharmacovigilence database in Section 6 were performed using our Python package available from github.com/bartroff792/mult-seq-dependence. First we describe the distributional set ups for our simulation studies, and in then discuss the results in Tables 2 and 3 and conclude by comparing our proposed sequential procedures with appropriate fixed-sample competitors.
@@ 6.1 Introduction
The United Kingdom’s Yellow Card Pharmacovigilence Database (yellowcard.mhra.gov.uk/information), run by the Medicines and Healthcare Products Regulatory Agency (MHRA), collects voluntary reports on safety and side-effects from the public and healthcare professionals in the UK on a host of healthcare treatments including medicines, vaccines, medical devices, and even e-cigarettes; in what follows, for simplicity we refer to the treatments in Yellow Card generically as “drugs.” The Yellow Card data collection scheme, which now includes a mobile phone app, began in 1964, spurred by the thalidomide crisis. Its use has grown steadily since then, now receiving more than 20,000 reports of possible side effects each year, and totaling more than half a million reports in the scheme’s first 40 years.
We created a Python script to download and analyze “Interactive Drug Analysis Profiles” PDF reports from Yellow Card on all of the roughly 2,800 different drugs in the database through February 2016. Figure 1 is a heatmap showing the number of drugs (indicated by color) in these reports as a function of their number of reported adverse action reactions ($y$ axis, on the $\log_{10}$ scale) and number of years ($x$ axis) the drug has been in the database, calculated from the starting date for collection of reports on the drug until the closure date of the drug’s most recent summary report. The figure shows that some of the drugs have many thousands of reaction reports, collected over decades. However, the majority of drug entries have fewer reports, collected over less than one decade, emphasizing the need for nimble, sequential monitoring of this data.
@@ 6.2 Illustrating the sequential procedures’ performance on the database
In order to generate a medically realistic correlation structure for the drugs, we utilized BioGPT (Luo et al.,, 2022), a generative language model trained on biomedical research articles. For each drug under consideration, we calculated the element-wise mean of the BioGPT’s token embedding for “[DrugName] as a drug” and then applied $K$-means clustering to obtain $K=15$ clusters of the drugs. Within each cluster, the copula method described in Section 5 was then used to generate the Poisson data using the Toeplitz covariance matrix given there, but with the correlation parameter $\rho$ generated as a Beta$(4,2)$ random variable, shifted and scaled to have support $[-1,1]$. In this way, drugs that were medically related according to BioGPT were more likely to receive non-trivial correlation.
In order to account accurately for the differing number of reports that drugs with different prescription and use rates in the database have, we set up the hypotheses tests not in terms of the Poisson rate parameters (36)-(37) but rather a scaled parameter
@@
The data generation scheme above is equivalent to a Binomial number of amnesia reports being generated at each time step, with success probability (38) and number of trials equal to the random total number of Poisson reports.
To set up our hypothesis tests, we utilize the database to calculate
@@
representing that the drug has a “typical” versus abnormally high rate of amnesia reports. For these we implemented the open-ended sequential test described in Theorem 3.1 for Binomial data using FDR controlled at $q_{1}=0.05$ and FNR at $q_{2}=0.15$ using with the Wald approximations (17)-(18) the BH step values (12). Running this data generation scheme with $1800$ data streams a single time is computationally intensive even on a Google Cloud Computing instance, so rather than report Monte Carlo statistics, in Table 4 we give the output of a typical run and there list the first 15 drugs whose streams terminated accept, and the first 15 to reject, the null hypothesis $H^{(j)}$. Since the simulation is set up where each “observation” is the number of reports in a year, in the table the Termination Step corresponds to the number of years required for termination of each stream. The Termination Level is a measure of the strength of the terminal action, smaller meaning a stronger acceptance/rejection of the null, and equals the cumulative value of the level $\ell$ in step 3a or 3b of the procedure as defined in Section 3.1.
While we draw no medical conclusions from our simulations, we note that some of the drugs the sequential procedure stopped to reject their null hypotheses in favor of an abnormally high level of amnesia reports, have been associated with amnesia in the biomedical literature (e.g., Chavant et al.,, 2011).

### v2

@@ Abstract
We propose sequential multiple testing procedures which control the false discovery rate (FDR) or the positive false discovery rate (pFDR) under arbitrary dependence between the data streams. This is accomplished by “optimizing” an upper bound on these error metrics for a class of step down sequential testing procedures. Both open-ended and truncated versions of these sequential procedures are given, both being able to control both the type I multiple testing metric (FDR or pFDR) at specified levels, and the former being able to control both the type I and type II (e.g., FDR and the false nondiscovery rate, FNR). In simulation studies, these procedures provide 45-65% savings in average sample size over their fixed-sample competitors. We illustrate our procedures on drug data from the United Kingdom’s Yellow Card Pharmacovigilance Database.
@@ 1 Introduction
Hart and Weiss, (1997) used linear programming to find a sharp upper bound on the probability of rejection for a class of multiple testing procedures in terms of the true, marginal c.d.f.s of the $p$-values. Calculating their bound therefore requires knowledge of at least the marginals of the true joint distribution. Guo and Rao, (2008) also used optimization but to find an upper bound on the false discovery rate (FDR) in terms of a step down procedure’s step values, that holds under arbitrary joint distribution of the $p$-values. By choosing the step values to control the upper bound, they produce a step down procedure whose FDR control holds under arbitrary joint distributions. The current paper extends this line of research in the following ways. First, by extending to the sequential setting an FDR controlling procedure without requiring or imposing assumptions on the joint distribution of the $p$-values; the open-ended versions of these sequential procedures allow the possibility of simultaneously controlling both type I and type II error metrics (e.g., FDR and the false nondiscovery rate, FNR). Secondly, by removing an implicit assumption on the $p$-values in Guo and Rao’s proof that is stronger than validity of $p$-values and applies to both the sequential and fixed-sample settings; see Appendix A for details. Finally, by extending this methodology to positive false discovery rate (pFDR) and its type II analog.
@@ 2.1 Data Streams, Hypotheses, and Error Metrics
In order to implement our proposed procedures, the marginal type I and II error probabilities of component test statistics on each stream will need to be controlled at certain levels; see (4)-(5) below. Beyond that, we make no assumptions about the dimension of the sequentially-observed data $X_{n}^{(j)}$, which may themselves be vectors of varying size, nor about the dependence structure of within-stream data $X_{n}^{(j)},X_{m}^{(j)}$ or between-stream data $X_{n}^{(j)},X_{m}^{(j^{\prime})}$ ($j\neq j^{\prime}$). In particular there can be arbitrary “overlap” between data streams, an extreme case being that all the data streams are the same, which is equivalent to testing multiple hypotheses about a single data source. For any positive integer $j$ let $[j]=\{1,\ldots,j\}$. For each data stream $j\in[J]$, assume that there is a parameter vector $\theta^{(j)}\in\Theta^{(j)}$ determining that distribution of the stream $X_{1}^{(j)},X_{2}^{(j)},\ldots$, and it is desired to test a null hypothesis $H^{(j)}$ versus an alternative hypothesis $G^{(j)}$, where $H^{(j)}$ and $G^{(j)}$ are disjoint subsets of the parameter space $\Theta^{(j)}$ containing $\theta^{(j)}$. It is not required that $H^{(j)}\cup G^{(j)}=\Theta^{(j)}$, e.g., one-sided alternatives or separated hypotheses are possible. The null hypothesis $H^{(j)}$ is considered true if $\theta^{(j)}\in H^{(j)}$, and false if $\theta^{(j)}\in G^{(j)}$. The global parameter $\theta=(\theta^{(1)},\ldots,\theta^{(J)})$ is the concatenation of the individual parameters and is contained in the global parameter space $\Theta=\Theta^{(1)}\times\cdots\times\Theta^{(J)}$.
@@
The general notation (1) includes fully-sequential sampling where the streamwise sample sizes may take any value $1,2,\ldots$ ad infinitum, but other sampling set ups fit this as well including group sequential, truncated, and even fixed sample size testing. For example, the $n$th “observation” $X_{n}^{(j)}$ in the $j$th stream may actually be the $n$th group $X_{n}^{(j)}=(X_{n,1}^{(j)},\ldots,X_{n,\ell}^{(j)})$ of size $\ell$. Moreover, the group size $\ell$ may vary with $n$ and may even be data-dependent, e.g., determined by some type of adaptive sampling. Similarly, truncated sequential (or group sequential) sampling can be implemented for the $j$th stream by defining $X_{n}^{(j)}=\emptyset$ for all $n>\overline{N}^{(j)}$ for some stream-specific truncation point $\overline{N}^{(j)}$, or globally for all streams by replacing statements like “for some $n$” in what follows with “for some $n\leq\overline{N}$,” for some global truncation point $\overline{N}$. One may represent a fixed sample size in the $j$th stream in this way by taking $\overline{N}^{(j)}=1$, or in all streams with $\overline{N}=1$.
@@
For any multiple testing procedure under consideration, let $V$ denote the number of true null hypotheses it rejects (i.e., the number of false positives), $W$ the number of false null hypotheses it accepts (i.e., the number of false negatives), and $R$ the number of null hypotheses it rejects. The number of null hypotheses accepted is therefore $J-R$. Under the true value of the parameter $\theta$, the false discovery and non-discovery rates (Benjamini and Hochberg,, 1995, FDR, FNR) are
@@
where $x\vee y=\max\{x,y\}$. Similarly, the positive false discovery rate pFDR (Storey,, 2002) and its type II analog, the positive false nondiscovery rate pFNR, are defined as
@@ 2.2 Test Statistics and Critical Values
The theorems in this paper will give upper bounds on the FDR and pFDR, and their type II versions, for sequential procedures implemented with arbitrary step values $\bm{\alpha},\bm{\beta}$. In practice we recommend using some of the commonly used values such as those recommended by Benjamini and Hochberg, (1995) or Benjamini and Liu, (1999). For desired value $q_{1}\in(0,1)$ of the type I error metric (FDR or pFDR), the former are
@@
Here $x\wedge y=\min\{x,y\}$. We recommend using the same expressions for $\bm{\beta}=\bm{\beta}(q_{2})$ for desired value $q_{2}\in(0,1)$ of the type II error metric (FNR or pFNR). In our examples below we will refer to (12) and (13) as the BH and BL step values, respectively.
@@ 2.3 Simple hypotheses and Wald approximations
In order to apply the above to the $j$th stream in the multiple testing set up for given step value vectors $\bm{\alpha}$ and $\bm{\beta}$, a slight extension of Bartroff and Song, (2020, Theorem 5.1) shows that choosing
@@
satisfies (4)-(5) up to Wald’s approximations, and that $\widetilde{\alpha}_{k}+\beta_{k}\leq 1$ and $\alpha_{k}+\widetilde{\beta}_{k}\leq 1$ for all $k\in[J]$. Although, strictly speaking, the inequalities (4)-(5) will only be guaranteed to hold approximately using these approximations, Hoel et al., (1971) show that the actual type I and II error probabilities can only exceed the nominal values by a negligibly small amount in the worst case, and the difference approaches 00 for small nominal values, which is relevant in the present multiple testing situation where we will utilize fractions of $\alpha$ and $\beta$. Alternatives to the Wald approximations in the simple vs. simple set up are Monte Carlo or to replace the terms in (16) by the values $\log\beta$ and $\log(1/\alpha)$, respectively, for which (14)-(15) hold conservatively (see Hoel et al.,, 1971) and proceed similarly.
@@ 4.3 Truncated, rejective procedures controlling FDR or pFDR
    regardless of the dependence between data streams, if $\gamma_{1}>0$ satisfies (25) for this procedure.
@@ 5 Simulation studies
In this section we present the results of simulation studies of the sequential step down procedures described above in Sections 3 and 4 in order to evaluate their operating characteristics and compare them to analogous fixed-sample procedures. All computations in this section and the analysis of the UK Yellow Card pharmacovigilance database in Section 6 were performed using our Python package available from github.com/bartroff792/mult-seq-dependence. First we describe the distributional set ups for our simulation studies, and in then discuss the results in Tables 2 and 3 and conclude by comparing our proposed sequential procedures with appropriate fixed-sample competitors.
@@ 6.1 Introduction
The United Kingdom’s Yellow Card pharmacovigilance Database (yellowcard.mhra.gov.uk/information), run by the Medicines and Healthcare Products Regulatory Agency (MHRA), collects voluntary reports on safety and side-effects from the public and healthcare professionals in the UK on a host of healthcare treatments including medicines, vaccines, medical devices, and even e-cigarettes; in what follows, for simplicity we refer to the treatments in Yellow Card generically as “drugs.” The Yellow Card data collection scheme, which now includes a mobile phone app, began in 1964, spurred by the thalidomide crisis. Its use has grown steadily since then, now receiving more than 20,000 reports of possible side effects each year, and totaling more than half a million reports in the scheme’s first 40 years.
We created a Python script to download and analyze “Interactive Drug Analysis Profiles” PDF reports from Yellow Card on all of the roughly 2,800 different drugs in the database through February 2016. Figure 1 is a heatmap showing the number of drugs (indicated by color) in these reports as a function of their total number of reported adverse action reactions over this period ($y$ axis, on the $\log_{10}$ scale) and number of years ($x$ axis) the drug has been in the database, calculated from the starting date for collection of reports on the drug until the closure date of the drug’s most recent summary report. The figure shows that some of the drugs have many thousands of reaction reports, collected over decades. However, the majority of drug entries have fewer reports, collected over less than one decade, emphasizing the need for nimble, sequential monitoring of this data.
@@ 6.2 Illustrating the sequential procedures’ performance on the database
In order to generate a medically realistic correlation structure for the drugs, we utilized BioGPT (Luo et al.,, 2022), a generative language model trained on biomedical research articles. For each drug under consideration, we calculated the element-wise mean of the BioGPT’s token embedding (its numerical representation of phrases, words, or characters) for “[DrugName] as a drug” and then applied $K$-means clustering to obtain $K=15$ clusters of the drugs. Within each cluster, the copula method described in Section 5 was then used to generate the Poisson data using the Toeplitz covariance matrix given there, but with the correlation parameter $\rho$ generated as a Beta$(4,2)$ random variable, shifted and scaled to have support $[-1,1]$. In this way, drugs that were medically related according to BioGPT were more likely to receive non-trivial correlation.
In order to account accurately for the differing number of reports that drugs with different prescription and use rates in the database have, we set up the hypothesis tests not in terms of the Poisson rate parameters (36)-(37) but rather a scaled parameter
@@
The data generation scheme above is equivalent to a Binomial number of amnesia reports being generated at each time step, with success probability (38) and number of trials equal to the Poisson random total number of reports.
To set up our hypothesis tests, we utilize the Yellow Card database to calculate
@@
representing that the drug has a “typical” versus abnormally high rate of amnesia reports. For these we implemented the open-ended sequential test described in Theorem 3.1 for Binomial data using FDR controlled at $q_{1}=0.05$ and FNR at $q_{2}=0.15$, using with the Wald approximations (17)-(18) the BH step values (12). Running this data generation scheme with $1800$ data streams a single time is computationally intensive even on a Google Cloud Computing instance, so rather than report Monte Carlo statistics, in Table 4 we give the output of a typical run and there list the first 15 drugs whose streams terminated to accept, and the first 15 to reject, the null hypothesis $H^{(j)}$. Since the simulation is set up where each “observation” is the number of reports in a year, in the table the Termination Step corresponds to the number of years required for termination of each stream. The Termination Level is a measure of the strength of the terminal action, where smaller values indicate a stronger acceptance/rejection of the null, and equals the cumulative value of the level $\ell$ in step 3a or 3b of the procedure as defined in Section 3.1.
6.3 Discussion
While we draw no medical conclusions from our simulations, we note that some of the drugs identified by our sequential procedure by stopping to reject their null hypotheses in favor of an abnormally high level of amnesia reports, have been associated with amnesia in the biomedical literature (e.g., Chavant et al.,, 2011). Monitoring pharmacovigilance databases for amnesia and other cognitive impairment adverse events is a real-world priority of statisticians and regulators around the world; see Borchert et al., (2019) and Sato et al., (2020) for examples of such analyses in the U.S.’s Food and Drug Administration Adverse Event Reporting System (FAERS; see www.fda.gov/media/97567/download), and Sunwoo et al., (2024) for an analysis of the Korean Adverse Event Reporting System Database. The example presented here provides a possible new statistical tool for this important task that takes into account the real-time, streaming nature of the data as well as the inherent multiple comparisons issues with such data. Although we are aware of proposed approaches that separately take into account the time-sequential nature of the data (e.g., Hocine et al.,, 2009) and the multiple comparison aspect, this is the first proposal to incorporate both that we are aware of.
The dominant statistical methods in this existing literature are “disproportionality analyses” (see Bihan et al.,, 2020; Lavertu et al.,, 2021; Zhang et al.,, 2024; Zorych et al.,, 2013), utilizing odds (or log odds) ratios of the probability of reporting an adverse event, which is equivalent to our use of the probabilities of such reports in (38)-(39). There we use the probabilities of such reports across the entire database as the levels of “typical” and “extreme” in (39) to set our hypotheses. This approach, known as “signal of disproportional reporting,” is common in the literature (e.g., Hauben et al.,, 2005; Hauben and Reich,, 2005; Zorych et al.,, 2013), although different approaches exist on how to choose these cutoffs and some of these authors use individual-level covariates, such as age in Hocine et al., (2009), in setting these cutoffs, which was not available in our example. However, signal of disproportional reporting is not a requirement of our method proposed here and those levels could be set by other methods such as from the medical literature, regulatory concerns, etc.
The approach proposed here could be used with other types of drugs and side effects, although domain-specific details the particular application will likely affect the choices made in implementation. For example, in monitoring adverse side effects of statins in the Korean pharmacovigilance database, Kim et al., (2017) grouped together statin-specific symptoms and measured severity on the WHO-specified scale, and the life-threatening nature of adverse immune-mediated reactions associated with certain immunotherapies (e.g., Ali and Watson,, 2017) would be taken into account in setting the early stopping properties of the procedure.